OpenStudy (anonymous):

A triangle has vertices at A(1, 5), B(-3, -3) & C(-3, 2). Classify the triangle by is sides.

OpenStudy (amistre64):

the base is 5; the sides are: 4^2 +8^2 = 16 +64 = sqrt(80) = 4 20 = 2 45 = 22 5 4sqrt(5) and the other is:

OpenStudy (amistre64):

more like this